OpenStudy (anonymous):

Solve using the quadratic formula x^2-7x-6=0

OpenStudy (anonymous):

Compare your quadratic equation with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\\\] \[ \\ \sqrt{b^2-4ac}=...?\] then the two roots of x are: \(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

I did that part.

OpenStudy (anonymous):

show us whatever you have got.

OpenStudy (anonymous):

and i simplified it and now im stuck on